A father is 3 times older than his son now. In 10 years time their combined ages will b 68. How old are they now?
Found 2 solutions by rwm, mananth:
Answer by rwm(914)

f=3s
(f+10)+(s+10)=68
f+s=48
4s=48
s=12
f=36

Answer by mananth(16946)

son =x
father = 3x
..
after 10 years
son = x+10
father= 3x+10
...
x+10+3x+10= 68
4x+20=68
4x=48
x=12 yera son's age
father = 36 years